OpenStudy (anonymous):

anyone good at Finite math? TI-84 calculator ?

OpenStudy (anonymous):

im good with Ti-84 but not at finite math though :DDD

OpenStudy (anonymous):

aw okay thank you. haha I just can;t figure out what to plug in for my problem

OpenStudy (anonymous):

A group of private investors purchased a condominium complex for $4 million. They made an initial down payment of 12% and obtained financing for the balance. If the loan is to be amortized over 12 years at an interest rate of 13% per year compounded quarterly, find the required quarterly payment.

OpenStudy (anonymous):

$ to be financed = 0.88*$4 millon = $3,520,000 r = 0.0325/qr (0.13/4), n = 48 qrs (12years*4 quarters) PMT = PVr/(1-(1+r)^-n) = $3,520,000*.0325/(1-1.0325^-48) = $114400/0.784584423=$145,809.6753 per quarter. Therefore the quarterly payment is $145,809.6753

OpenStudy (anonymous):

you are awesome thanks!!!!

OpenStudy (anonymous):

you're welcome :DDD

OpenStudy (anonymous):

you're very welcome